# Computation and Communication Co-scheduling for Timely Multi-Task Inference at the Wireless Edge

### v1

@@ 3       Initialize $\pi_{m,j}(t)\leftarrow 0$ for all $(m,j)$
4       Get $\boldsymbol{\lambda}^{*}$ and $\mu^{*}$ that maximizes $\bar{p}(\boldsymbol{\lambda},\boldsymbol{\mu};t:T)$
@@ 14            $A(t)=A(t)\setminus(m^{*},j^{*})$
Algorithm 1 Maximum Gain First (MGF) Policy
@@ VI Numerical Experiments
(a) Dis. Sum of Errors vs. $rk_{m}$
(b) Dis. Sum of Errors vs. $N$
(c) Dis. Sum of Errors vs. $M$
Figure 3: Model-based evaluations: (a) discounted sum of errors vs. number of tasks per source $rk_{m}$, (b) discounted sum of errors vs. number of channels $N$, and (c) discounted sum of errors vs. number of sources $M$. The MGF policy has the potential to achieve a discounted sum of inference errors up to 26 and 32 times lower than the MAF and Random policies, respectively.
@@
  • Maximum Gain First (MGF) Policy: The scheduling decision under this policy follows Algorithm 1.
@@
We evaluate these three policies under two scenarios:
@@
  • Real-world evaluations: We conduct two machine learning experiments and incorporate the resulting inference error functions into the simulation (Sec. VI-B).
@@ VI-A Synthetic Evaluations
In this section, we use three AoI penalty functions: $p_{m,j}(\delta)=\delta,\exp(0.5\delta),10\text{log}(\delta)$. Each function is assigned to one-third of the inference tasks in each source $m$.
@@
Fig. 3(a) illustrates the discounted sum of inference errors versus the number of tasks per source ($rk_{m}$) over a time horizon of $T=100$. Referring to Definition 2, we set $k_{m}=3$ and vary $r$. The additional simulation parameters are $M=20$, $N=10$, $\gamma=0.9$, $n_{m,j}=1$ for all tasks $(m,j)$, $C_{m}=2$ for all sources $m$, and $w_{m,j}=0.01$ for half of the tasks and $1$ for the other half. We see that, when $rk_{m}=15$, the total discounted penalty for the MAF policy is $26$x higher than that of the MGF policy, while the random policy incurs $32$x higher penalty. The performance of the MAF policy deteriorates more rapidly than the MGF policy as the number of tasks per source increases, aligning with our findings in Theorem 1.
@@
In Fig. 3(b), we plot the discounted sum of errors against the number of channels $N$. Here, $k_{m}=3$ and $r=3$ for each source $m$, and the rest of the parameters are the same as in Fig. 3(a). We see that increasing $N$ improves performance for all policies, but more rapidly for MGF. When $N=2$, the MAF policy incurs four times penalty of the MGF policy. This performance gap narrows as $N$ increases, but even with $N=20$, the MAF policy’s inference error remains twice that of the MGF policy.
@@
In Fig. 3(c), we plot the discounted sum of errors against the number of sources $M$, with other parameters the same as in Fig. 3(a)&(b). As the number of sources increases, we see that the performance gap between MAF and MGF policies widens. This shows that MGF is more effective as the number of sources competing for MTRI resources increases.
(a) Inference Error vs AoI
(b) Dis. Sum of Errors vs. $N$
(c) Dis. Sum of Errors vs. $M$
Figure 4: Data-driven evaluations: (a) Inference error vs. AoI for segmentation task (left), where we use 100 (1-IoU) loss function, and for traffic prediction (right), where we use quadratic loss function. (b) Discounted sum of errors vs. number of channels $N$ in which two inference tasks are given priority with weight $1$ and the others have weight $0.01$. (c) Discounted sum of errors vs. number of sources $M$ in which half of the inference tasks given priority with weight $1$ and the others have weight $0.01$. MGF outperforms the other two policies.
VI-B Real-World Evaluations
VI-B1 Machine Learning Experiments
@@
Hidden Layers 16
Input and Output Size 1
@@
Fig. 4(a) illustrates the resulting inference errors vs. AoI.
VI-B2 Evaluations of Scheduling Policies
We now evaluate the scheduling policies employing these inference error functions. In Fig. 4(b), we plot the discounted sum of errors against the number of channels $N$ over a time horizon of $T=100$, with two inference tasks $k_{m}=2$ per source and scaling factor $r=1$. We set $M=20$, $\gamma=0.9$, $n_{m,j}=1$, and $C_{m}=2$ for all sources. Task weights $w_{m,j}$ are set to 1 for tasks (1,2) and (5,1), and 0.01 for the rest. As expected, increasing $N$ is seen to improve performance across all policies. Notably, when $N=4$, the MGF policy outperforms the MAF policy by $10\%$. Additionally, Fig. 4(b) clearly demonstrates the consistently poor performance of the random policy.
Fig. 4(c) illustrates the performance of the scheduling policies as the number of sources $M$ increases, over a finite horizon of $T=100$. Each source has two inference tasks $k_{m}=2$ and $r=1$, with other simulation parameters set to $N=10$, $\gamma=0.9$, $n_{m,j}=1$, and $C_{m}=1$. We assign weights $w_{m,j}$ of $1$ to half the inference tasks and $0.01$ to the rest. We see that while the MAF and MGF policies perform similarly with a small number of sources, the MGF policy becomes better as the number of sources increases.
VI-C Discussion
Synthetic evaluations (Sec. VI-A) demonstrate that the MGF policy significantly outperforms both the MAF and random policies, achieving up to 26x and 32x better performance, respectively. The improvement is particularly evident as the number of tasks per source increases, thus validating Theorem 1. Real-world data-driven evaluations (Sec. VI-B) further confirm the MGF policy’s superiority, though the gains are not as high as the gains achieved in synthetic evaluations. This is due to limitations in the inference tasks, specifically, the small number of tasks per source (only 2) and the small variation in the inference error functions as AoI increases (Fig. 4(a)). This limited variation is attributed to the dataset’s unchanging background scene and consistent number of vehicles over time. However, with more tasks or greater variability in inference difficulty, MGF’s advantage would be expected to increase. In scenarios where feature freshness impacts inference difficulty, MGF’s adaptability will be especially beneficial. Under such conditions, MGF’s performance gains should approach those seen in synthetic evaluations. Overall, the simulation results highlight strong performance of the MGF policy compared to baselines.
@@ VII Conclusion
In this paper, we studied the problem co-scheduling computation and communication in MTRI systems to minimize inference errors under resource constraints. We formulated this problem as a weakly-coupled MDP with inference errors described as penalty functions of AoI. To address the resulting PSPACE-hard complexity, we developed a novel MGF policy, which our theoretical analysis proved to be asymptotically optimal as the number of inference tasks increases. Numerical evaluations using both synthetic and real-world datasets further validated MGF’s superior performance compared to baseline policies.
@@ Appendix A Proof of Lemma 2
In this proof, we use $n_{m,j}=1$ for the simplicity of analysis. Denote $\pi^{\text{MGF}}_{m,j}(t)$ as the action under MGF policy. Define
@@
$C^{*}_{m}(t)=\sum_{j=1}^{k_{m}}\pi^{*}_{m,j}(t),N^{*}(t)=\sum_{m=1}^{M}\sum_{j=1}^{k_{m}}\pi^{*}_{m,j}(t)$ (22)
Case 1: At time $t$, all constraints are satisfied under policy $\pi^{*}$. In this case, we have $|I_{t}|=0.$ Case 2: At least one constraint does not satisfy under policy $\pi^{*}$. In this case, if any sub-problem $(m,j)\in I_{t}$, then $\pi^{*}_{m,j}(t)=1$ and $\pi^{\text{MGF}}_{m,j}(t)=0$ due to resource limitation, i.e., $C^{*}_{m}(t)>C_{m}$ or $N^{*}(t)>N$ or both. Because the active action $\pi^{*}_{m,j}(t)=1$ consumes one communication and one computation resource, we have
@@
$\mathbb{E}[(C^{*}_{m}(t)-C_{m})^{+}]^{2}\overset{(a)}{\leq}\mathbb{E}[(C^{*}_{m}(t)-\mathbb{E}[C^{*}_{m}(t)])^{+}]^{2}$
$\!\!\!\!\!\!\!\overset{(b)}{\leq}\text{Var}(C^{*}_{m}(t))\overset{(c)}{\leq}\sum_{j=1}^{k_{m}}\bigg{(}E[\pi^{*}_{m,j}(t)]-\mathbb{E}[\pi^{*}_{m,j}(t)]^{2}\bigg{)}{\leq}k_{m},$ (24)
where (a) holds because on average $\mathbb{E}[C^{*}_{m}(t)]\leq C_{m}$, see [12, Proposition 3.2(c)], (b) holds due to Jensen’s inequality, (c) is because of Bhatia-Davis inequality. Similarly, we can have
$\mathbb{E}[(N^{*}(t)-N)^{+}]^{2}\leq\text{Var}(N^{*}(t))\leq\sum_{m=1}^{M}k_{m}.$ (25)
By taking an average on (23) and substituting (A) and (25) into (23), we obtain
$\mathbb{E}[|I_{t}|]\leq\sum_{m=1}^{M}\sqrt{k_{m}}+\sqrt{\sum_{m=1}^{M}k_{m}}.$ (26)

### v2

@@ 3   Initialize $\pi_{m,j}(t)\leftarrow 0$ for all $(m,j)$
4 Get $\boldsymbol{\lambda}^{*}$ and $\mu^{*}$ that maximizes $\bar{p}(\boldsymbol{\lambda}(t),\boldsymbol{\mu}(t);t:T)$
@@ 15    $A(t)=A(t)\setminus(m^{*},j^{*})$
Algorithm 1 Reoptimized Maximum Gain First Policy
Algorithm 1 presents our reoptimized maximum gain first (MGF) scheduler for solving the main problem (4)-(6). We denote $\pi_{\text{MGF}}$ by the policy provided in Algorithm 1. At each time $t$, the policy $\pi_{\text{MGF}}$ prioritizes generating and transmitting features ($\pi_{m,j}(t)=1$) for the inference tasks with highest gain index, while adhering to the available communication and computation resources. Our policy then proceeds as follows:
  • (1)
    First, our policy re-optimizes $\boldsymbol{\lambda}^{*}(t)$ and $\mu^{*}(t)$ by maximizing $\bar{p}(\boldsymbol{\lambda}(t),\boldsymbol{\mu}(t);t:T)$. Then, the gain indices $\alpha_{m,j,t}(\Delta_{m,j}(t))$ for all tasks $(m,j)$ defined in (1) are calculated.
  • (2)
    Let $A(t)$ be the set of inference tasks with positive gain indices:
    $A(t)=\bigg{\{}(m,j):\alpha_{m,j,t}(\Delta_{m,j}(t))>0\bigg{\}}$ (22)
    Then, our policy selects the inference task $(m^{*},j^{*})$ that satisfies
  • (3)
V-B Performance Analysis
We now analyze the performance of our policy relative to the original problem (4)-(6). Following standard practice in the weakly-coupled MDP literature [13, 45], a set of sub-problems at source $m$ are said to be in the same class if they share identical penalty functions, weights, and transition probabilities.
Definition 2 (Asymptotic optimality).
Consider a “base” MTRI system with $N$ channels, $M$ sources, $k_{m}$ classes of sub-problems per source $m$, a computation resource budget $C_{m}$ for source $m$ and $N$ shared communication resources by $M$ sources. Now, consider a “multiplied system”, where each class of sub-problems per source $m$ contains $r$ inference tasks, each source $m$ has $rC_{m}$ computation resources and $rN$ communication resources, while maintaining a constant $M$ sources. Let $\bar{p}^{r}_{{\text{MGF}}}$ and $\bar{p}^{r}_{{\pi}_{opt}}$ represent the discounted infinite horizon sum of inference errors under policy $\pi_{\text{MGF}}$ and an optimal policy for the multiplied system, respectively. The policy $\pi_{\text{MGF}}$ is asymptotically optimal if $\bar{p}^{r}_{{\text{MGF}}}=\bar{p}^{r}_{{\pi}_{opt}}$ for all $\pi\in\Pi$ as inference task per class $r$ approaches $\infty$, i.e.,
$\lim_{r\to\infty}\bar{p}^{r}_{{\text{MGF}}}=\bar{p}^{r}_{{\pi}_{opt}}.$ (24)
First, we establish Lemma 2, which is a key tool to showing asymptotic optimality. We define a policy $\pi^{*}=(\pi^{*}_{m,j}(t))_{\forall m,j,t\leq T}$, where $\pi^{*}_{m,j}(t)$ is the reoptimized action obtained by using Lemma 1 with the Lagrange multipliers that maximize $\bar{p}_{t}(\boldsymbol{\lambda}_{t:T},\boldsymbol{\mu}_{t:T})$ defined in (11). Using (13) and (1), we can verify that $\pi^{*}_{m,j}(t)=1$ if $\alpha_{m,j,t}(\Delta_{m,j}(t))>0$.
Lemma 2.
For any time $t$ and AoI values, the expected number of subproblems with actions that differ between the reoptimized MGF policy provided in Algorithm 1 and the policy $\pi^{*}$ is bounded from above by
$\sum_{m=1}^{M}\sqrt{k_{m}}+\sqrt{\sum_{m=1}^{M}k_{m}}.$
Proof.
See Appendix A. ∎
Then, we can obtain our main theoretical result:
Theorem 1.
If there exists a finite constants $\bar{p}_{l}$ and $\bar{p}_{h}$ such that $\bar{p}_{l}\leq w_{m,j}p_{m,j}(\delta)\leq\bar{p}_{h}$ for any sub-problem $(m,j)$ and AoI value $\delta=1,2,\ldots,$ and
$T\geq\log_{\frac{1}{\gamma}}\left(\sum_{m=1}^{M}\sqrt{rk_{m}}\right),$ (25)
then the MGF policy is asymptotically optimal as the number of inference tasks $r$ for all classes $(m,j)$ increases to infinite. Specifically, we have
$=O\bigg{(}\frac{1}{\sum_{m=1}^{M}\sqrt{rk_{m}}}\bigg{)},$ (26)
where $k_{m}$ is the number of sub-problems per source, $M$ is the number of sources, and $\gamma$ is the discount factor.
Proof.
See Appendix B. ∎
While prior work has introduced gain-index-based policies for RMAB problems [36, 42, 11], these cannot be directly applied to general weakly-coupled MDPs. Our gain-index-based policy, a specialized re-optimized fluid policy (Definition 3) [13], achieves tighter asymptotic optimality for MTRI systems (see (1)) compared to the $O(\frac{1}{\sqrt{\sum_{m=1}^{M}rk_{m}}})$ bound in [13]. This improvement is obtained by using Lemma 2, which strengthens the result [13, Lemma EC1.1] by exploiting the MTRI constraint structure: sub-problems utilize only their source’s computational resources. Unlike the general system in [13] where all resources are globally shared, MTRI systems have local (computational) and global (communication) resources, yielding a tighter bound. For example, a sub-problem associated with source $m_{1}$ would only consume computational resources from source $m_{1}$ and not from another source $m_{2}$, unlike in [13], where all sub-problems share all resources in the system.
1 for $t=0,1,\ldots$ do
2 Update $\Delta_{m,j}(t)$ for all $(m,j)$
3   Initialize $\pi_{m,j}(t)\leftarrow 0$ for all $(m,j)$
4 Get $\lambda_{1}^{*},\lambda^{*}_{2},\ldots,\lambda^{*}_{M}$, and $\mu^{*}$ that maximizes $\bar{p}_{t}(\lambda_{1},\lambda_{2},\ldots,\lambda_{M},\mu)$
5 $\alpha_{m,j}\leftarrow\alpha_{m,j,t}(\Delta_{m,j}(t))$ for all $(m,j)$
6
7 $C_{m,\text{curr}}\leftarrow 0$ and $N_{\text{curr}}\leftarrow 0$
8 $A(t)\leftarrow\{(m,j):\alpha_{m,j}>0)\}$
9 while $A(t)$ is not empty do
10    $(m^{*},j^{*})\leftarrow\operatorname*{arg\,max}_{m,j}\alpha_{m,j}$
12    if $c\leq C_{m^{*}}$ and $n\leq N$ then
13        Update $\pi_{m^{*},j^{*}}(t)\leftarrow 1$
14         Update $C_{m^{*},\text{curr}}\leftarrow c$ and $N_{\text{curr}}\leftarrow n$
15    $A(t)=A(t)\setminus(m^{*},j^{*})$
Algorithm 2 Simplified Reoptimized MGF Policy
V-C Simplified Reoptimized Maximum Gain First Policy
In Algorithm 1, at every time $t$, we require to optimize the Lagrange variables $\boldsymbol{\lambda}_{t:T}$ and $\boldsymbol{\mu}_{t:T}$. The definition (8)-(10) show that the total number of parameters to optimize is (M+1)(Tt)(M+1)(T-t) at each time $t$. The computational complexity of such optimization may not be feasible for real-time implementation. To handle this issue, we keep the Lagrange variables $\lambda_{m,t}=\lambda_{m}$ and $\mu_{t}=\mu$ for all $t=1,2,\ldots,T$. This reduces the number of optimization variables to $(M+1)$. After adopting time-invariant Lagrange variables, the primal problem (IV-A) becomes:
$\bar{p}_{\tau}(\lambda_{1},\lambda_{2},\ldots,\lambda_{M},\mu)=$
$\;\inf_{\pi\in\Pi}\sum_{t=\tau}^{T}\sum_{m=1}^{M}\sum_{j=1}^{k_{m}}\frac{\gamma^{t-\tau}\mathbb{E}_{\pi}\left[w_{m,j}p_{m,j}(\Delta_{m,j}(t))\right]}{K}$
$+\sum_{m=1}^{M}\lambda_{m}\sum_{t=\tau}^{T}\frac{\gamma^{t-\tau}}{K}\left(\left(\sum_{j=1}^{k_{m}}\pi_{m,j}(t)\right)-C_{m}\right)$
$+{\mu}\sum_{t=\tau}^{T}\frac{\gamma^{t-\tau}}{K}\left(\left(\sum_{m=1}^{M}\sum_{j=1}^{k_{m}}\pi_{m,j}(t)n_{m,j}\right)-N\right),$ (27)
and the dual problem (11) becomes
$(\lambda_{1}^{*},\lambda^{*}_{2},\ldots,\lambda^{*}_{M},\mu^{*})=\operatorname*{arg\,max}_{\lambda_{1},\lambda_{2},\ldots,\lambda_{M},\mu\geq 0}\bar{p}(\boldsymbol{\lambda}(\tau),\boldsymbol{\mu}(\tau);\tau:T).$ (28)
The dual sub-gradient ascent algorithm for solving (28) is as follows:
$\lambda_{m}(i+1)=\max\bigg{\{}\lambda_{m}(i)+$
$\frac{\beta_{m}}{Ki}\left(\sum_{t=\tau}^{T}\sum_{j=1}^{k_{m}}\gamma^{t-\tau}\pi_{m,j}^{*}(t)-\frac{(1-\gamma^{(T-t)})C_{m}}{(1-\gamma)}\right),0\bigg{\}},$ (29)
$\mu(i+1)=\max\bigg{\{}\mu(i)+$
$\frac{\beta}{Ki}\left(\sum_{t=\tau}^{T}\sum_{m=1}^{M}\sum_{j=1}^{k_{m}}\gamma^{t-\tau}\pi_{m,j}^{*}(t)n_{m,j}-\frac{(1-\gamma^{(T-t)})N}{(1-\gamma)}\right),0\bigg{\}},$ (30)
where $\beta,\beta_{m}>0$ are the step sizes and $\pi_{m,j}^{*}(t)$ is the optimal action to the $(m,j)$-th sub-problem with $\lambda_{m}(i)$ and $\mu(i)$ obtained in the $i$-th iteration.
@@ VI Numerical Experiments
Figure 3: Dis. Sum of Errors vs. $rk_{m}$
Figure 4: Dis. Sum of Errors vs. NN
Figure 5: Dis. Sum of Errors vs. $M$
@@
  • Maximum Gain First (MGF) Policy: For numerical study, we use the modified Algorithm 2.
@@
We evaluate these three policies under three scenarios:
@@
  • Remote Robot Car Detection: We conduct a remote robot car detection experiment. We generate the inference error function for robot car detection and incorporate the resulting inference error functions into the simulation (Sec. VI-B)
  • Traffic Prediction and Segmentation: In this experiment, we consider two machine learning tasks: (i) scene segmentation and (ii) traffic prediction on the NGSIM dataset [53, 54, 55, 56]. Then, we incorporate the resulting inference error functions into the simulation (Sec. VI-C).
@@ VI-A Synthetic Evaluations
In this section, we use three AoI penalty functions: $p_{m,j}(\delta)=\delta,\exp(0.5\delta),10\text{log}(\delta)$. These functions are widely used in AoI literature as estimation error [14, 15]. Each function is assigned to one-third of the inference tasks in each source $m$.
@@
Fig. 3 illustrates the discounted sum of inference errors versus the number of tasks per source ($rk_{m}$) over a time horizon of $T=100$. Referring to Definition 2, we set $k_{m}=3$ and vary $r$. The additional simulation parameters are $M=20$, $N=10$, $\gamma=0.9$, $n_{m,j}=1$ for all tasks $(m,j)$, $C_{m}=2$ for all sources $m$, and $w_{m,j}=0.01$ for half of the tasks and $1$ for the other half. We see that, when $rk_{m}=15$, the total discounted penalty for the MAF policy is 2626x higher than that of the MGF policy, while the random policy incurs 3232x higher penalty. The performance of the MAF policy deteriorates more rapidly than the MGF policy as the number of tasks per source increases, aligning with our findings in Theorem 1.
@@
In Fig. 4, we plot the discounted sum of errors against the number of channels $N$. Here, $k_{m}=3$ and $r=3$ for each source $m$, and the rest of the parameters are the same as in Fig. 4. We see that increasing $N$ improves performance for all policies, but more rapidly for MGF. When $N=2$, the MAF policy incurs four times penalty of the MGF policy. This performance gap narrows as $N$ increases, but even with $N=20$, the MAF policy’s inference error remains twice that of the MGF policy.
@@
In Fig. 5, we plot the discounted sum of errors against the number of sources $M$, with other parameters the same as in Fig. 3&4. As the number of sources increases, we see that the performance gap between MAF and MGF policies widens. This shows that MGF is more effective as the number of sources competing for MTRI resources increases.
Figure 6: System Model of Experiment
Figure 7: Inference Error ($1-\mathrm{IoU}$) vs. AoI
Figure 8: Dis. Sum of Errors vs. NN
Figure 9: Dis. Sum of Errors vs. NN
Figure 10: One sample of prediction results for robot car 33 with different AoI values
Figure 11: One sample of prediction results for robot car 22 with different AoI values
VI-B Remote Robot Car Detection
In this section, we discuss the system model and the results on remote robot car detection experiment.
Fig. 6 illustrates the experimental system model, which consists of 4 sources and 5 robot cars. Source 1 observes 2 of the cars, while the remaining 3 sources each monitor a single car. Each source uses an onboard camera and the YOLO11x[57] model to detect robots within its view. The model generates bounding box information for each detection. In each time slot, the sources transmit the information of the detection to a receiver. However, the system is subject to two constraints: (i) A maximum of $N$ bounding boxes can be transmitted across the network and (ii) Source $1$ can only detect and transmit information for one of its two observed cars. The receiver’s prediction for a car’s location is the most recently delivered bounding box information for that car.
Fig. 7 plots the inference error, defined as $1-\mathrm{IoU}$, versus the AoI for the $5$ robot cars. The Intersection over Union (IoU) measures the overlap between the predicted and ground-truth bounding boxes. To measure the inference error of a robot car, we have used $800$ data samples for each AoI value. One sample of prediction results for AoI values =2,5,10,15,20=2,5,10,15,20 are illustrated in Fig. 10 and Fig. 11, respectively for robot car $3$ and $2$. The error for robot cars 2 and 4 rapidly approaches its maximum value of 1, because these cars move much faster than the others. We can observe from Figs 10&11 that robot car 2 is moving faster compared to robot car $3$ which yields larger inference error for robot car 2 compared to robot car 3.
Fig. 8 plots the discounted sum of the normalized inference error over $T=100$ time slots against the number of channels, $N$. For this simulation, we set the weight $w_{1,2}=5$, other weights are set to $1$, and discount factor $\gamma=0.9$. Our proposed “Reoptimized MGF” policy clearly outperforms the “MAF” and “Random” baselines. The Random policy performs the worst as it does not use any state information. While the MAF policy considers the AoI, it neglects the system dynamics—specifically, how AoI impacts inference error. In contrast, our Reoptimized MGF policy leverages this system dynamics, leading to its superior performance.
Fig. 9 plots the discounted sum of the normalized inference error over $T=100$ time slots against the weight, $w_{1,2}$ set for robot car $2$ observed by source $1$. For this simulation, other weights are set to $1$, and discount factor is set to $\gamma=0.1$. This plot illustrates that as the weight $w_{1,2}$ increases, the performance gap between our policy and other baselines increases.
VI-C Traffic Prediction and Segmentation
Figure 12: Inference Error vs AoI
Figure 13: Dis. Sum of Errors vs. NN
Figure 14: Dis. Sum of Errors vs. $M$
@@ VI-C Traffic Prediction and Segmentation
Hidden units per layer 16
Input and output dimensions 1
@@
Fig. 12 illustrates the resulting inference errors vs. AoI.
We now evaluate the scheduling policies employing these inference error functions. In Fig. 13, we plot the discounted sum of errors against the number of channels $N$ over a time horizon of $T=100$, with two inference tasks $k_{m}=2$ per source and scaling factor $r=1$. We set $M=20$, $\gamma=0.9$, $n_{m,j}=1$, and $C_{m}=2$ for all sources. Task weights $w_{m,j}$ are set to 1 for tasks (1,2) and (5,1), and 0.01 for the rest. As expected, increasing $N$ is seen to improve performance across all policies. Notably, when $N=4$, the MGF policy outperforms the MAF policy by $10\%$. Additionally, Fig. 13) clearly demonstrates the consistently poor performance of the random policy.
Fig. 14 illustrates the performance of the scheduling policies as the number of sources $M$ increases, over a finite horizon of $T=100$. Each source has two inference tasks $k_{m}=2$ and $r=1$, with other simulation parameters set to $N=10$, $\gamma=0.9$, $n_{m,j}=1$, and $C_{m}=1$. We assign weights $w_{m,j}$ of $1$ to half the inference tasks and $0.01$ to the rest. We see that while the MAF and MGF policies perform similarly with a small number of sources, the MGF policy becomes better as the number of sources increases.
@@ VII Conclusion
In this paper, we studied the computation and communication co-scheduling problem in MTRI systems to minimize inference errors under resource constraints. We formulated this problem as a weakly-coupled MDP with inference errors described as penalty functions of AoI. To address the resulting PSPACE-hard complexity, we developed a novel reoptimized MGF policy, which our theoretical analysis proved to be asymptotically optimal as the number of inference tasks increases. We also discussed how to simplify a reoptimized policy by reducing the number of optimization variables. Numerical evaluations using both synthetic and real-world datasets further validated our reoptimized MGF’s superior performance compared to baseline policies. For synthetic evaluations, we use three different types of inference error functions that are widely used in AoI literature. For real-world experiments, remote robot car detection and vehicular inference tasks are studied.
@@ Appendix A Proof of Lemma 2
Firstly, we denote the action at time $t$ for $j$-th inference task of $m$-th source under the reoptimized MGF policy provided in Algorithm 1 by $\pi^{\text{MGF}}_{m,j}(t)$. We denote the reoptimized action at time $t$ for $j$-th inference task of $m$-th source under the policy $\pi^{*}$ by $\pi^{*}_{m,j}(t)$, where $\pi^{*}$ maximizes $\bar{p}_{t}(\boldsymbol{\lambda}_{t:T},\boldsymbol{\mu}_{t:T})$ defined in (11).
Next, we denote the number of subproblems with actions that differ between the reoptimized MGF policy provided and the policy $\pi^{*}$ at time $t$ by $I_{t}$:
@@
$C^{*}_{m}(t)=\sum_{j=1}^{k_{m}}\pi^{*}_{m,j}(t),N^{*}(t)=\sum_{m=1}^{M}\sum_{j=1}^{k_{m}}\pi^{*}_{m,j}(t),$ (31)
where we use $n_{m,j}=1$ for the simplicity of analysis in this proof.
Case 1: At time $t$, all constraints are satisfied under policy $\pi^{*}$. In this case, we have $|I_{t}|=0.$ Case 2: At least one constraint does not satisfy under policy $\pi^{*}$. In this case, if a sub-problem $(m,j)\in I_{t}$, then $\pi^{*}_{m,j}(t)=1$ and $\pi^{\text{MGF}}_{m,j}(t)=0$ due to resource limitation, i.e., $C^{*}_{m}(t)>C_{m}$ or $N^{*}(t)>N$ or both. Because the active action $\pi^{*}_{m,j}(t)=1$ consumes one communication and one computation resource, we can upper bound $I_{t}$ by
@@
$\mathbb{E}[(C^{*}_{m}(t)-C_{m})^{+}]^{2}$
$\overset{(a)}{\leq}\mathbb{E}[(C^{*}_{m}(t)-\mathbb{E}[C^{*}_{m}(t)])^{+}]^{2}$
$\overset{(b)}{\leq}\text{Var}(C^{*}_{m}(t))$
$\overset{(c)}{\leq}\sum_{j=1}^{k_{m}}\bigg{(}E[\pi^{*}_{m,j}(t)]-\mathbb{E}[\pi^{*}_{m,j}(t)]^{2}\bigg{)}$
${\leq}k_{m},$ (33)
where (a) holds because on average $\mathbb{E}[C^{*}_{m}(t)]\leq C_{m}$, see [13, Proposition 3.2(c)], (b) holds due to Jensen’s inequality, (c) is because of Bhatia-Davis inequality. Similarly, we can have
$\mathbb{E}[(N^{*}(t)-N)^{+}]^{2}$
$\leq\mathbb{E}[(N^{*}(t)-\mathbb{E}[N^{*}])^{+}]^{2}$
$\leq\text{Var}(N^{*}(t))$
$\leq\sum_{m=1}^{M}k_{m}.$ (34)
By taking an average on (32) and substituting (A) and (A) into (32), we obtain
$\mathbb{E}[|I_{t}|]$ $=\sum_{m=1}^{M}\mathbb{E}[(C^{*}_{m}(t)-C_{m})^{+}]+\mathbb{E}[(N^{*}(t)-N)^{+}]$
$\leq\sum_{m=1}^{M}\sqrt{k_{m}}+\sqrt{\sum_{m=1}^{M}k_{m}}.$ (35)
This concludes the proof.